The Downtown Community Barbecue served 238 dinners. A child’s plate cost $3.10 and an adult’s plate cost $8.10. A total of $1,297.80 was collected. How many of each type of plate was served?
Round answers to the nearest whole person.


Answer by rfer(16322)

c+a=238
a=238-c
-------------
3.10c+8.10a=1297.80
3.10c+8.10(238-c)=1297.80
3.10c+1927.80-8.10c=1297.80
-5c=1297.80-1927.80
-5c=-630
c=-630/-5
c=126 children
238-126=112 adults